OpenStudy (anonymous):

2. Factor completely 2x²y+12xy+18y

OpenStudy (anonymous):

What can you pull out of each one of them?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

and if you pull out a 2, what do you get when you factor?

OpenStudy (anonymous):

2x²y+12xy+18y =2y (x²+6x+9) ...

OpenStudy (anonymous):

oh, makes sense...

OpenStudy (anonymous):

@matricked is right because you can pull out 2y and not just 2

OpenStudy (anonymous):

you want to find a common factor(s) in all of them

OpenStudy (anonymous):

and then you have a quadratic

OpenStudy (anonymous):

=2y(x²+2*3x+9) ..[ (a+b)²= a² +2ab +b²) ] =2y(x+3)²

OpenStudy (anonymous):

confused.. haha.

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

all of it. i used an online factoring calculator and the answer said 6y(2x²+2x+3) but i have to show work and idk i'm just lost

OpenStudy (anonymous):

i'm awful at math.. haha.

OpenStudy (anonymous):

=2y(x²+2*3x+9) =2y(x² +3x +3x +9) =2y( x(x+3) +3(x+3) ) =2y(x+3)(x+3) =2y(x+3)²

OpenStudy (anonymous):

sooo.. is 2y(x+3)² the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, thank you

OpenStudy (anonymous):

check your question first..

OpenStudy (anonymous):

6y(2x²+2x+3) = 12 x²y +12xy +18 y..

OpenStudy (anonymous):

soooo...... that's the answer?

OpenStudy (anonymous):

but you have posted 2 x²y +12xy +18 y..

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ooooookay soooooooo is the answer 2y(x+3)² or 6y(2x²+2x+3)

OpenStudy (anonymous):

the question is asking you to factor completely

OpenStudy (anonymous):

so, factoring completely would be 2y(x+3)²

OpenStudy (anonymous):

it would be the 12x^2y...

OpenStudy (anonymous):

if that is what the question wants

OpenStudy (anonymous):

can you check your answer?

OpenStudy (anonymous):

so wait 12x²y is the answer?

OpenStudy (anonymous):

do you have the answer?

OpenStudy (anonymous):

uh the answer online said 6y(2x²+2x+3) but you said something different and the other person said the same thing as the online calculator said but then he confused me and now i'm just all freaking confused.

OpenStudy (anonymous):

wait is the answer 2y(x+3)(x-3)?

OpenStudy (anonymous):

2y(x+3)² factors to 2x^2+12xy+18y and that is what you started with

OpenStudy (anonymous):

so, it's right?

OpenStudy (anonymous):

2y(x+3)(x+3) is how I would leave your answer

OpenStudy (anonymous):

alright, thanks

OpenStudy (anonymous):

2x²y+12xy+18y = 2y(x+3)(x+3) but for 12 x²y +12xy +18 y = 6y(2x²+2x+3)

OpenStudy (anonymous):

you just typed the same equation with different answers..?

OpenStudy (anonymous):

matricked is showing the online calculator that you used solved for a different polynomial than you may have intended

OpenStudy (anonymous):

notice the first term is 12 for the online calculator and for your actual answer it has a 2

OpenStudy (anonymous):

ohhhhhhhhhhhhh